Recent Posts

Pages: [1] 2 3 ... 10
1
501 Single Host / Re: orientation of inductor coil
« Last post by JPK on Today at 05:37:44 PM »
Yes, correct.
2
CP4500 Stereo Bus Compressor / Re: Make up gain no longer working
« Last post by JPK on Today at 05:34:43 PM »
Check if you get the same voltage variation on TP12 (0 to 2V3) with MIX at 100% and no input.
3
LA502T Tube Optical Compressor / Re: test procedure - having trouble
« Last post by JPK on Today at 05:26:43 PM »
The values you measured on TP8 and U3-U4 are correct so you should be able to adjust the LEDs.
One possible cause of the problem could be an incorrect value for one of the resistors R29 to R33. Please double-check their values.
4
501 Single Host / orientation of inductor coil
« Last post by skrzyp on November 02, 2026, 06:12:50 PM »
To be safe, just checking on this before soldering. Is the orientation pictured correct? My part is different than the one pictured in the build guide. Thanks in advance.
5
CP4500 Stereo Bus Compressor / Re: Make up gain no longer working
« Last post by MRMNW on October 28, 2026, 01:41:54 PM »
Anything else I can check?
6
LA502T Tube Optical Compressor / Re: test procedure - having trouble
« Last post by doktar on October 27, 2026, 08:05:58 PM »
Replaced the blown LED's.  

Repeated testing, the GR LEDS are still over what they should be, and can't adjust lower because all the trimmers are all ready at the lowest position.

Any advice or should I just live with it?

7
CP4500 Stereo Bus Compressor / Re: Make up gain no longer working
« Last post by MRMNW on October 27, 2026, 10:45:20 AM »
Thank you

Yes to both of these
8
CP4500 Stereo Bus Compressor / Re: Make up gain no longer working
« Last post by JPK on October 26, 2026, 05:24:59 PM »
Check if the center pin of P2 (makeup) goes from 0 to +12VDC when moving.
Next check that this voltage reaches R90 and U22 pin7 (with less amplitude 0-2V4).
9
CP4500 Stereo Bus Compressor / Make up gain no longer working
« Last post by MRMNW on October 26, 2026, 01:29:32 PM »
Hi,
The unit has been working fine up until recently, now the makeup gain does nothing. 

I'm getting a reading of -1.42v at TP12 as per step 1.2 on the voltage checks and it doesn't change when moving the makeup gain on step 1.3.

Any help appreciated

Thanks
10
LA502T Tube Optical Compressor / Re: test procedure - having trouble
« Last post by doktar on October 23, 2026, 06:34:07 PM »
I believe I've blown out the LED's, when testing on the IC's I accidentally shorted pins with the meter probe,  I've ordered some replacements and will report back.
Pages: [1] 2 3 ... 10
SMF spam blocked by CleanTalk